### v1

# A theorem in relation to quantum Ising models

Abstract
The physics of the paradigmatic one-dimensional transverse field quantum Ising model $J\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+h\sum_{i}\sigma^{z}_{i}$ is well-known. Instead, let us imagine “applying” the transverse field via a transverse Ising coupling of the spins to partner auxiliary spins, i.e. $H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\text{partner of }i}$. If each spin of the chain has a unique auxiliary partner, then the theorem states that the resultant eigenspectrum is still the same as that of the quantum Ising model with $h=J_{z}$ and the degeneracy of the entire spectrum is $2^{\text{number of auxiliary spins}}$. This follows from the existence of extensively large and mutually anticommuting sets of conserved quantities for $H$. We can interpret this situation as the auxiliary spins remaining paramagnetic down to zero temperature with an extensive ground state degeneracy. This is lost upon the loss of the unique partner condition for the full spin chain. Other cases where such extensive degeneracy survives are also discussed.
Setup: Consider the Hamiltonian
$H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\partial i}$ (1)
where $\partial i$ stands for the auxiliary partner of site $i$ on the spin chain. Several examples are seen in the figures that follow.
(a) Refer to caption
(b) Refer to caption      (c) Refer to caption
(d) Refer to caption      (e) Refer to caption
Figure 1: Examples of quantum Ising chains with different configurations for the auxiliary spins.
Case (a) corresponds to when all sites on the chain obey the unique auxiliary partner condition. Case (b-e) corresponds to when not all sites on the chain obey the unique auxiliary partner condition. Case (d-e) corresponds to when all sites on the chain violate the unique auxiliary partner condition.
In all cases, we have the standard global $Z_{2}$ symmetry of the tranverse field quantum Ising model (TFQIM). It may be implemented as a $180^{\circ}$ rotation around the $z$-axis, i.e.
$\mathcal{U}^{Z_{2}}=\prod_{i}\otimes\;\mathcal{R}^{\pi,z}_{i}$ (2)
with
$\mathcal{R}^{\pi,z}_{i}=e^{i\pi\sigma^{z}_{i}}.$ (3)
Under this
$\sigma^{z}_{i}$ $\rightarrow\mathcal{U}^{Z_{2}}\sigma^{z}_{i}{\mathcal{U}^{Z_{2}}}^{\dagger}=\mathcal{R}^{\pi,z}_{i}\sigma^{z}_{i}\mathcal{R}^{-\pi,z}_{i}=\sigma^{z}_{i}$ (4)
$\sigma^{x}_{i}$ $\rightarrow-\sigma^{x}_{i}$ (5)
and $\text{and }H$ $\rightarrow H$ (6)
The consequence of this is the conservation of the parity of total chain magnetization in the $z$-direction $M^{z}=\sum_{i}\sigma^{z}_{i}$. Clearly, $[H,M^{z}]\neq 0$ but
$[H,M^{z}\text{ mod }2]$ $=\sum_{\langle i,j\rangle}[\sigma^{x}_{i}\sigma^{x}_{j},M^{z}\text{ mod }2]=0$ (7)
This conservation is also equivalent to fermion parity conservation after the Jordan-Wigner transformation which maps the Ising term $\sigma^{x}_{i}\sigma^{x}_{j}$ to a sum of hopping and superconducting terms in the fermion language. Also in all cases, $\sigma^{z}_{\partial i}$ is conserved for all $i$, i.e.
$[H,\sigma^{z}_{\partial i}]=0$ (8)
as can be verified easily. This in fact facilitates the computation of the exact eigenspectrum via the Jordan-Wigner transformation [1, 2].
Let us start with case (a) in Fig. 1 which satisfies the unique auxiliary partner condition for all sites of the chain.
Theorem: For this case, the theorem states
  • The resultant eigenspectrum is still the same as that of the quantum Ising model with $h=J_{z}$.
  • The degeneracy of the spectrum is $2^{N_{\partial i}}$ where $N_{\partial i}$ is the number of the auxiliary spins.
Interpretation:
  • The above can be interpreted as the auxiliary spins remaining paramagnetic down to zero temperature co-existing with Ising order/disorder on the chain  [2].
  • This also implies an extensive ground state degeneracy.
To prove the theorem we use the following lemma.
Lemma: Let there be two conserved quantities $A$ and $B$, i.e. $[H,A]=[H,B]=0$, that are mutually anticommuting $\{A,B\}=0$. For an eigenstate in the $A$-basis, i.e. $H|\psi\rangle=E|\psi\rangle$ and $A|\psi\rangle=a|\psi\rangle$, there exists another state $|B\psi\rangle\equiv B|\psi\rangle$ which is also an eigenstate with $H|B\psi\rangle=E|B\psi\rangle$ and $A|B\psi\rangle=-a|B\psi\rangle$. If $A$ has no zero eigenvalues, then $|B\psi\rangle$ is distinct than $|\psi\rangle$.
Proof: There are additional conserved quantities which are absent in the TFQIM. These are $\sigma^{x}_{i}\sigma^{x}_{\partial i}$, i.e.
$[\sigma^{x}_{i}\sigma^{x}_{\partial i},H]=0$ (9)
for all $i$ as can be verified easily. Furthermore these conserved quantities anticommute with $\sigma^{z}_{i}$, i.e.
$\{\sigma^{x}_{i}\sigma^{x}_{\partial i},\sigma^{z}_{\partial i}\}=0$ (10)
for all $i$ as can be verified easily. Also both sets of conserved quantities square to non-zero values and thus have no zero eigenvalues
$\left(\sigma^{z}_{i}\right)^{2}$ $=1$ (11)
$\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)^{2}$ $=1$ (12)
Thus by the application of the lemma above, for each eigenstate $|\psi\rangle$ of $H$, one arrives at $N_{\partial i}$ degenerate eigenstates as $\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)|\psi\rangle$. In fact there are many more degenerate eigenstates arrived at by the operation of the product of $\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)$ over any subset of the auxiliary partner sites. One can convince oneself that the total degeneracy is thus $2^{N_{\partial i}}$.
The eigenspectrum is same as that of TFQIM with $h=J_{z}$ is seen by choosing that sector of the Hamiltonian which corresponds to all the conserved $\sigma^{z}_{\partial i}$ being all up or all down, i.e. $\prod_{\partial i}\otimes|\uparrow^{z}_{\partial i}\rangle$ or $\prod_{\partial i}\otimes|\downarrow^{z}_{\partial i}\rangle$.
Now let us consider the case (b) in Fig. 1. Here again we have the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ with unique partners. For the two sites which share a partner, the conserved quantity is now $\sigma^{x}_{i}\sigma^{x}_{\partial(i,i+1)}\sigma^{x}_{i+1}$. This also anticommutes with $\sigma^{z}_{\partial(i,i+1)}$. Thus we can make similar arguments as above. In case (c), the conserved quantity is $\sigma^{x}_{i}\sigma^{x}_{\partial(i,i+1)}\sigma^{x}_{i+1}\sigma^{x}_{\partial(i+1,i+2)}\sigma^{x}_{i+2}$ with similar physics since in all the above cases (a-c) there are an extensive number of additional conserved quantities.
In case (d), the unique partner condition is lost for the full spin chain. Thus in this case we do not have an extensive number of additional conserved quantities. There is only one such quantity, i.e. $\prod_{i,\partial(i,i+1)}\otimes\sigma^{x}_{i}\otimes\sigma^{x}_{\partial(i,i+1)}$. This will lead to a degeneracy of 2 of the spectrum. The configuration of the auxiliary spins which corresponds to the ground state sector also needs determination [2]. Case (e) is another such example to show why the physics present in cases (a-c) is absent in the TFQIM.
Discussion: Let us consider case (a) for this discussion. It is natural to block diagonalize the Hamiltonian $H$ in terms of the conserved spin configurations of the auxiliary partner spins $\prod_{\partial i}\otimes|\sigma^{z}_{\partial i}\rangle$. However, the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ begs the following question: How to understand the physics if we were to organize the Hamiltonian blocks in terms of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ Firstly, fixing the configuration of the auxiliary spins as $\prod_{\partial i}\otimes|\sigma^{z}_{\partial i}\rangle$ implies no fluctuation in them. But fixing the eigenvalues (of ±$\pm 1$) of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ does not imply any such thing. In this way of block diagonalization, both the spins of the spin chain and the auxiliary spins keep fluctuating. This suggests that the (local) conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ has a gauge like character. From this point of view, for a given eigenstate $|\psi\rangle$, we can obtain degenerate eigenstates as $\sigma^{z}_{\partial i}|\psi\rangle$ or $\prod_{\{\partial j\}\subseteq\{\partial i\}}\sigma^{z}_{\partial j}|\psi\rangle$ for any subset of auxiliary spins. This again gives a degeneracy of $2^{N_{\partial i}}$ as expected. Due to this extensive degeneracy, the gauge charges or eigenvalues of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ can also keep fluctuating. This is because we can linearly combine the eigenstates from different gauge charge sectors to obtain a new eigenstate. Under time evolution, this linear combination will stay put, i.e. both the gauge charges and the auxiliary spin states keep fluctuating for all times.
Another perspective is to look at the same physics after the Jordan-Wigner transformation. Then we arrive at
$H=J_{x}\sum_{\langle i,j\rangle}\left(c^{\dagger}_{i}c_{j}+c^{\dagger}_{i}c^{\dagger}_{j}+\text{h.c. }\right)+J_{z}\sum_{i}(2n_{i}-1)(2n_{\partial i}-1)$ (13)
The global fermion parity is again conserved due to global $Z_{2}$ symmetry, but we also have local $Z_{2}$ symmetries in terms of local $180^{\circ}$ rotations around the $x$-axis for site $i$ and $\partial i$ which keep the Hamiltonian unchanged. This implies the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ on $(i,\partial i)$ bonds. Upon Jordan-Wigner transformation, we get
[…]
However, there are no kinetic hopping or superconducting terms ($^{\dagger}_{i}c_{\partial i}+c^{\dagger}_{i}c^{\dagger}_{\partial i}+\text{h.c.}\right)$.c. ) corresponding to the local $Z_{2}$ charges on $(i,\partial i)$ bonds. Thus all gauge sectors are degenerate. Also the mutual anticommutation of
$\{c^{\dagger}_{i}c_{\partial i}+c^{\dagger}_{i}c^{\dagger}_{\partial i}+\text{h.c.},n_{\partial i}\}=0$ (15)
implies that local $Z_{2}$ charges can fluctuate along with $n_{\partial i}$.
In conclusion, this note describes a mechanism for gauge like behaviour in quantum spin-$\frac{1}{2}$ systems. For any Hamiltonian, if it hosts mutually anticommuting sets of conserved quantities that have extensive cardinality, such behaviour would manifest. This may be a novel mechanism for the emergence of gauge-like physical degrees of freedom, e.g. when comparing to the Levin-Wen model, Kitaev’s honeycomb model, Haah’s code and X-cube model  [3, 4, 5, 6, 7]. This mechanism can in general operate in any number of dimensions. Constructing solvable models in two and higher dimensions based on this mechanism will be interesting. Some physical consequences originating from this mechanism has been discussed in Ref. [2].
Acknowledgements: Funding support from SERB-DST, India via Grant No. MTR/2022/000386 and partially by Grant No. CRG/2021/003024 is acknowledged.