### v3

# A theorem on extensive ground state entropy and some related models

Abstract
The physics of the paradigmatic one-dimensional transverse field quantum Ising model $J\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+h\sum_{i}\sigma^{z}_{i}$ is well-known. Instead, let us imagine “applying” the transverse field via a transverse Ising coupling of the spins to partner auxiliary spins, i.e. $H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\text{partner of }i}$. If each spin of the chain has a unique auxiliary partner, then the theorem states that the resultant eigenspectrum is still the same as that of the quantum Ising model with $\frac{h}{J}=\frac{J_{z}}{J_{x}}$ and the degeneracy of the entire spectrum is $2^{\text{number of auxiliary spins}}$. This follows from the existence of extensively large and mutually anticommuting sets of conserved quantities for $H$. We can interpret this as the auxiliary spins remaining paramagnetic down to zero temperature and an extensive ground state entropy. This is lost upon the loss of the unique partner condition for the full spin chain. Other cases where such degeneracy survives or gets lost are also discussed. Thus this theorem forms the basis for an exact statement on the existence of extensive ground state entropy.
Exact statements are of immense value in quantum many-body physics. They include exactly solvable models of course, but also go beyond them. Well-known examples of the second kind are Peierl’s argument for classical Ising models [1, 2] and Elitzur’s theorem in the context of lattice gauge theories that forbids local orders [3] with implications for the spontaneous symmetry breaking in superconductors [4]. Other semi-rigorous to rigorous examples are the Ginzburg criterion on the validity of mean-field theories [5], and Harris and Imry-Ma criteria on the effect of disorder on in clean systems [6, 7, 8, 9]. In this work, we will make an exact statement of this second kind and illustrate it through both various models including solvable ones. The statement concerns a mechanism that forces an extensive ground state entropy on a system as we shall see. The physics of systems with extensive ground state entropy are often interesting, e.g. classical spin ices [10] and the SYK model [11]. These should not be thought of as violating the third law of thermodynamics [12], but rather as capturing the relevant physics of real systems operating at the relevant temperature scales in a physical setting or experiment [13]. With this motivation, we go towards the theorem starting with its relation to quantum Ising models.
Consider the Hamiltonian
$H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\partial i}$ (1)
where $\partial i$ stands for the auxiliary partner of site $i$ on the spin chain. Several examples are seen in Fig. 1. We will stick to ferromagnetic couplings throughout in this article without loss of generality.
(a) Refer to caption
(b) Refer to caption      (c) Refer to caption
(d) Refer to caption      (e) Refer to caption
Figure 1: Examples of quantum Ising chains with different configurations for the auxiliary spins.
Case (a) corresponds to when all sites on the chain obey the unique auxiliary partner condition mentioned in the abstract. Case (b-e) corresponds to when not all sites on the chain obey the unique auxiliary partner condition. Case (d-e) corresponds to when all sites on the chain violate the unique auxiliary partner condition.
In all cases, we have the standard global $Z_{2}$ symmetry of the tranverse field quantum Ising model (TFQIM). It may be implemented as a $180^{\circ}$ rotation around the $z$-axis, i.e.
$\mathcal{U}^{Z_{2}}=\prod_{i}\otimes\;\mathcal{R}^{\pi,z}_{i}$ (2)
with
$\mathcal{R}^{\pi,z}_{i}=e^{i\pi\sigma^{z}_{i}/2}.$ (3)
Under this
$\sigma^{z}_{i}$ $\rightarrow\mathcal{U}^{Z_{2}}\sigma^{z}_{i}{\mathcal{U}^{Z_{2}}}^{\dagger}=\mathcal{R}^{\pi,z}_{i}\sigma^{z}_{i}\mathcal{R}^{-\pi,z}_{i}=\sigma^{z}_{i}$ (4)
$\sigma^{x}_{i}$ $\rightarrow-\sigma^{x}_{i}$ (5)
and $\text{and }H$ $\rightarrow H$ (6)
The consequence of this is the conservation of the parity of total chain magnetization in the $z$-direction $M^{z}=\sum_{i}\sigma^{z}_{i}$. Clearly, $[H,M^{z}]\neq 0$ but
$[H,M^{z}\text{ mod }2]$ $=\sum_{\langle i,j\rangle}[\sigma^{x}_{i}\sigma^{x}_{j},M^{z}\text{ mod }2]=0$ (7)
This conservation is also equivalent to fermion parity conservation after the Jordan-Wigner transformation [14] which maps the Ising term $\sigma^{x}_{i}\sigma^{x}_{j}$ to a sum of hopping and superconducting terms in the fermion language. Also in all cases, $\sigma^{z}_{\partial i}$ is conserved for all $i$, i.e.
$[H,\sigma^{z}_{\partial i}]=0$ (8)
as can be verified easily. This in fact facilitates the computation of the exact eigenspectrum via the Jordan-Wigner transformation [15].
Let us start with case (a) in Fig. 1 which satisfies the unique auxiliary partner condition for all sites of the chain. For this case, we have the following
Theorem:–
  • The resultant eigenspectrum is still the same as that of the quantum Ising model with $\frac{h}{J}=\frac{J_{z}}{J_{x}}$.
  • The degeneracy of the spectrum is $2^{N_{\partial i}}$ where $N_{\partial i}$ is the number of the auxiliary spins.
Interpretation:–
  • The above can be interpreted as the auxiliary spins remaining paramagnetic down to zero temperature co-existing with Ising order/disorder on the chain  [15].
  • This also implies an exponentially large ground state degeneracy in system size or extensive ground state entropy.
  • One may call this ground state as a co-existence state of Ising order/disorder with a “classical” spin liquid.
To prove the theorem we use the following lemma.
Lemma:– Let there be two conserved quantities $A$ and $B$, i.e. $[H,A]=[H,B]=0$, that are mutually anticommuting $\{A,B\}=0$. For an eigenstate in the $A$-basis, i.e. $H|\psi\rangle=E|\psi\rangle$ and $A|\psi\rangle=a|\psi\rangle$, there exists another state $|B\psi\rangle\equiv B|\psi\rangle$ which is also an eigenstate with $H|B\psi\rangle=E|B\psi\rangle$ and $A|B\psi\rangle=-a|B\psi\rangle$. If $A$ has no zero eigenvalues, then $|B\psi\rangle$ is distinct than $|\psi\rangle$.
Proof:– There are additional conserved quantities which are absent in the TFQIM. These are $\sigma^{x}_{i}\sigma^{x}_{\partial i}$, i.e.
$[\sigma^{x}_{i}\sigma^{x}_{\partial i},H]=0$ (9)
for all $i$ as can be verified easily. Furthermore these conserved quantities anticommute with $\sigma^{z}_{i}$, i.e.
$\{\sigma^{x}_{i}\sigma^{x}_{\partial i},\sigma^{z}_{\partial i}\}=0$ (10)
for all $i$ as can be verified easily. Also both sets of conserved quantities square to non-zero values and thus have no zero eigenvalues
$\left(\sigma^{z}_{i}\right)^{2}$ $=1$ (11)
$\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)^{2}$ $=1$ (12)
Thus by the application of the lemma above, for each eigenstate $|\psi\rangle$ of $H$, one arrives at $N_{\partial i}$ degenerate eigenstates as $\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)|\psi\rangle$. In fact there are many more degenerate eigenstates arrived at by the operation of the product of $\left(\sigma^{x}_{i}\sigma^{x}_{\partial i}\right)$ over any subset of the auxiliary partner sites. One can convince oneself that the total degeneracy is thus $2^{N_{\partial i}}$. The eigenspectrum is same as that of TFQIM with $\frac{h}{J}=\frac{J_{z}}{J_{x}}$ is seen by choosing that sector of the Hamiltonian which corresponds to all the conserved $\sigma^{z}_{\partial i}$ being all up or all down, i.e. $\prod_{\partial i}\otimes|\uparrow^{z}_{\partial i}\rangle$ or $\prod_{\partial i}\otimes|\downarrow^{z}_{\partial i}\rangle$.
Now let us consider the case (b) in Fig. 1. Here again we have the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ with unique partners. For the two sites which share a partner, the conserved quantity is now $\sigma^{x}_{i}\sigma^{x}_{\partial(i,i+1)}\sigma^{x}_{i+1}$. This also anticommutes with $\sigma^{z}_{\partial(i,i+1)}$. Thus we can make similar arguments as above. In case (c), the conserved quantity is $\sigma^{x}_{i}\sigma^{x}_{\partial(i,i+1)}\sigma^{x}_{i+1}\sigma^{x}_{\partial(i+1,i+2)}\sigma^{x}_{i+2}$ with similar physics since in all the above cases (a-c) there are an extensive number of additional conserved quantities.
In case (d), the unique partner condition is lost for the full spin chain. Thus in this case we do not have an extensive number of additional conserved quantities. There is only one such quantity, i.e. $\prod_{i,\partial(i,i+1)}\otimes\sigma^{x}_{i}\otimes\sigma^{x}_{\partial(i,i+1)}$. This will lead to a degeneracy of 2 of the spectrum. The configuration of the auxiliary spins which corresponds to the ground state sector also needs determination [15]. Case (e) is another such example to show why the physics present in cases (a-c) is absent in the TFQIM.
Discussion:– Let us consider case (a) for this discussion. It is natural to block diagonalize the Hamiltonian $H$ in terms of the conserved spin configurations of the auxiliary partner spins $\prod_{\partial i}\otimes|\sigma^{z}_{\partial i}\rangle$. However, the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ begs the following question: How to understand the physics if we were to organize the Hamiltonian blocks in terms of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ Firstly, fixing the configuration of the auxiliary spins as $\prod_{\partial i}\otimes|\sigma^{z}_{\partial i}\rangle$ implies no fluctuation in them. But fixing the eigenvalues (of ±$\pm 1$) of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ for all $i$ does not imply any such thing. In this way of block diagonalization, both the spins of the spin chain and the auxiliary spins keep fluctuating. This suggests that the (local) conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ has a gauge like character. From this point of view, for a given eigenstate $|\psi\rangle$, we can obtain degenerate eigenstates as $\sigma^{z}_{\partial i}|\psi\rangle$ or $\prod_{\{\partial j\}\subseteq\{\partial i\}}\sigma^{z}_{\partial j}|\psi\rangle$ for any subset of auxiliary spins. This again gives a degeneracy of $2^{N_{\partial i}}$ as expected. Due to this extensive degeneracy, the gauge charges or eigenvalues of the conserved $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ can also keep fluctuating. This is because we can linearly combine the eigenstates from different gauge charge sectors to obtain a new eigenstate. Under time evolution, this linear combination will stay put, i.e. both the gauge charges and the auxiliary spin states keep fluctuating for all times.
Another perspective is to look at the same physics after the Jordan-Wigner transformation. Then we arrive at
$H=J_{x}\sum_{\langle i,j\rangle}\left(c^{\dagger}_{i}c_{j}+c^{\dagger}_{i}c^{\dagger}_{j}+\text{h.c. }\right)+J_{z}\sum_{i}(2n_{i}-1)(2n_{\partial i}-1)$ (13)
The global fermion parity is again conserved due to global $Z_{2}$ symmetry, but we also have local $Z_{2}$ symmetries in terms of local $180^{\circ}$ rotations around the $x$-axis for site $i$ and $\partial i$ which keep the Hamiltonian unchanged. This implies the conservation of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ on $(i,\partial i)$ bonds. Upon Jordan-Wigner transformation, we get
[…]
However, there are no kinetic hopping or superconducting terms ($^{\dagger}_{i}c_{\partial i}+c^{\dagger}_{i}c^{\dagger}_{\partial i}+\text{h.c.}\right)$.c. ) corresponding to the local $Z_{2}$ charges on $(i,\partial i)$ bonds. Thus all gauge sectors are degenerate. Also the mutual anticommutation of
$\{c^{\dagger}_{i}c_{\partial i}+c^{\dagger}_{i}c^{\dagger}_{\partial i}+\text{h.c.},n_{\partial i}\}=0$ (15)
implies that local $Z_{2}$ charges can fluctuate along with $n_{\partial i}$.
Extensions:– Let us continue with case (a). Since $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ is conserved, the following Hamiltonian
$H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\partial i}+J^{\prime}_{x}\sum_{i}\sigma^{x}_{i}\sigma^{x}_{\partial i}$ (16)
also is solvable. However $\sigma^{z}_{\partial i}$ is not conserved anymore. Thus the extensive degeneracy will be lost. The spectrum now will depend on the conserved value of $\sigma^{x}_{i}\sigma^{x}_{\partial i}$ on all $(i,\partial i)$ bonds. E.g. the ground state will correspond to $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=1$ for $J^{\prime}_{x}>0$. Correponding to $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=1$, there are two states |±$|\pm^{x}_{i}\pm^{x}_{\partial i}\rangle$ on $(i,\partial i)$ bond. The $\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}$ term will keep $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=1$ unchanged. $\sigma^{z}_{i}\sigma^{z}_{\partial i}$ will flip between the two states |±$|\pm^{x}_{i}\pm^{x}_{\partial i}\rangle$ on $(i,\partial i)$ bond. Thus the above reduces to an effective (dual) TFQIM once the value of $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ is chosen on all $(i,\partial i)$ bonds. We may write it as follows
$H=J^{\text{eff}}\sum_{\langle i,j\rangle}\tau^{z}_{i}\tau^{z}_{j}+h^{\text{eff}}\sum_{i}\tau^{x}_{i}+J^{\prime}_{x}\sum_{i}\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ (17)
where the $\tau$ operators operate on the two states consistent with $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$, and $J^{\text{eff}}\propto J_{x}$, $h^{\text{eff}}\propto J_{z}$. One will again obtain the TFQIM spectrum in any conserved sector. The loss of extensive degeneracy corresponding to $J^{\prime}_{x}=0$ is seen through the third term above $J^{\prime}_{x}\sum_{i}\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$. One sees that there are still degenerate excited sectors given by different configurations of $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ which keep the sum $\sum_{i}\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ fixed. The degeneracies are basically $N_{i}\text{ choose }N_{\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=1}$. These degeneracies will be further broken down in presence of additional terms where solvability is likely not possible. By a similar token, for the following Hamiltonian
$H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}+J^{\prime}_{z}\sum_{\langle i,j\rangle}\sigma^{z}_{i}\sigma^{z}_{j}+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\partial i}$ (18)
$\sigma^{x}_{i}\sigma^{x}_{\partial i}$ is not conserved anymore, but $\sigma^{z}_{\partial i}$ stays conserved. Thus the extensive degeneracy will again be lost.
Figure 2: Ladder geometry with bond-dependent couplings as discussed in the text.
However, solving for the spectrum using the Jordan-Wigner transformation is not that straightforward.
Also may be noted that the following ladder Hamiltonian
$H=J_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}$ $+J_{\partial x}\sum_{\langle i,j\rangle}\sigma^{x}_{\partial i}\sigma^{x}_{\partial j}$
$+J_{z}\sum_{i}\sigma^{z}_{i}\sigma^{z}_{\partial i}+J^{\prime}_{x}\sum_{i}\sigma^{x}_{i}\sigma^{x}_{\partial i}$ (19)
as shown in Fig. 2 is effectively equivalent to
$H=\sum_{\langle i,j\rangle}J^{\text{eff}}_{ij}\tau^{z}_{i}\tau^{z}_{j}+h^{\text{eff}}\sum_{i}\tau^{x}_{i}+J^{\prime}_{x}\sum_{i}\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ (20)
with $h^{\text{eff}}\propto J_{z}$. The case of $J^{\text{eff}}_{ij}$ requires more attention. For the (ground state) sector corresponding to $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=1$ on all $(i,\partial i)$ bonds, $J^{\text{eff}}_{ij}\propto(J_{x}+J_{\partial x})$ independent of the bond location. The same would be true for the sector corresponding to $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=-1$ on all $(i,\partial i)$ bonds. Recall we are considering ferromagnetic couplings in this article throughout. For other sectors where $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle$ is not uniformly the same sign, the bond location becomes important. For a bond $(i,j)$ such that $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle=\langle\sigma^{x}_{j}\sigma^{x}_{\partial j}\rangle$, $J^{\text{eff}}_{ij}\propto(J_{x}+J_{\partial x})$. For a bond $(i,j)$ such that $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle\neq\langle\sigma^{x}_{j}\sigma^{x}_{\partial j}\rangle$, $J^{\text{eff}}_{ij}$ itself fluctuates between $\propto(J_{x}-J_{\partial x})$ and $\propto-(J_{x}-J_{\partial x})$ depending on the state of the spins on the $(i,\partial i)$ and $(j,\partial j)$ bonds. Obtaining the spectrum in these excited sectors is therefore more involved. For $J_{x}=J_{\partial x}$ which would be the case in presence of mirror symmetry between the two legs of the ladder, there occurs a simplication and $J^{\text{eff}}_{ij}=0$ on those bonds where $\langle\sigma^{x}_{i}\sigma^{x}_{\partial i}\rangle\neq\langle\sigma^{x}_{j}\sigma^{x}_{\partial j}\rangle$. This leads to disconnected TFQIM segments which can again be solved for the exicted eigenspectrum. The above ladder Hamiltonian can be looked at as a solvable local spinless fermionic model for $J^{\prime}_{x}=0$,
[…]
$-t_{\partial x}\sum_{\langle i,j\rangle}(c^{\dagger}_{\partial i}c_{\partial j}+c^{\dagger}_{\partial i}c^{\dagger}_{\partial j}+\text{ h.c.})+V\sum_{i}n_{i}n_{\partial i}$
The physical situation is that of two $p$-wave superconducting wires coupled through a short-ranged Couloumbic interaction. Analogous physics will carry through in this context. We may conjecture that the physics extends to situation when the hopping and pairing amplitudes are not exactly equal. A stronger conjecture would be the stability of the ground state in presence of hopping and/or superconducting amplitudes between the two wires. The ground state of the fermionic system will be two locked superconducting ground states independent of $V$. The solvable case of $J^{\prime}_{x}\neq 0$ leads to non-local terms in the fermionic situation and may be ignored.
Till now the discussion has been limited to one dimensional systems. Let us now consider a two-dimensional spin model where the mechanism underlying the theorem – existence of mutually anticommuting sets of extensively large conserved quantities – is operational. Consider the bond-dependent Hamiltonian sketched in Fig. 3.
Figure 3: A two-dimensional spin model with extensive ground state degeneracy.
The system is composed of a) square plaquettes with either $\sigma^{x}_{i}\sigma^{x}_{j}$ or $\sigma^{z}_{i}\sigma^{z}_{j}$ couplings exclusively, b) crosses composed of both $\sigma^{x}_{i}\sigma^{x}_{j}$ and $\sigma^{z}_{i}\sigma^{z}_{j}$ segments criss-crossing each other, and c) hexagonal plaquettes with alternating $\sigma^{x}_{i}\sigma^{x}_{j}$ and $\sigma^{z}_{i}\sigma^{z}_{j}$ couplings as a result of a) and b).
Let us first look at the conserved quantites. They are
  1. 1.
    $\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k}\sigma^{z}_{l}$ on the “$\sigma^{x}_{i}\sigma^{x}_{j}$” square plaquettes.
  2. 2.
    $\sigma^{x}_{i}\sigma^{x}_{j}\sigma^{x}_{k}\sigma^{x}_{l}$ on the “$\sigma^{z}_{i}\sigma^{z}_{j}$” square plaquettes.
  3. 3.
    $\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k}$ on the “$\sigma^{x}_{i}\sigma^{x}_{j}$” segments of the crosses.
  4. 4.
    $\sigma^{x}_{i}\sigma^{x}_{j}\sigma^{x}_{k}$ on the “$\sigma^{z}_{i}\sigma^{z}_{j}$” segments of the crosses.
The conserved nature of these quantities may be verified easily. There does not seem to be any obvious conserved quantity associated with the hexagonal plaquettes. All the above quantities form extensively large sets. The first and second sets of conserved quantities commute with each other. The third and fourths sets anticommute with each other. Similarly, the first and fourth sets anticommute with each other and the second and third sets anticommute with each other. Here, by anticommute, we mean those quantities that have common sites between them. The first and third sets commute with each other, and the second and fourth sets commute with each other.
Eigenspectrum solvability of this model is not apparent, but by the application of the lemma, we can conclude that this system will also host an extensive ground state entropy. In fact, the full eigenspectrum will be massively degenerate in this sense. The counting can be ascertained by first spanning all the sites with mutually conserved sets from the above, and then counting the remaining sets that anticommute (in the sense of this paper) with the spanning sets. The maximum of all possible ways of doing this will give the entropy due to this mechanism. In this model, if we use first and second sets as the spanning sets, then the remaining sets contribute an entropy of $4\ln 2\;k_{B}$ per unit cell. Instead, if we use first and third sets as the spanning sets, then the remaining sets contribute an entropy of $3\ln 2\;k_{B}$ per unit cell. Similary, if we use second and fourth sets as the spanning sets, then the remaining sets again contribute an entropy of $3\ln 2\;k_{B}$ per unit cell. Thus the ground state entropy is $4\ln 2\;k_{B}$ per unit cell through this mechanism. This model and its natural deformations perhaps deserves further study as also hinted below.
Further discussion:– This work describes a way to construct spin models with extensive ground state entropy. For any Hamiltonian, if it hosts mutually anticommuting sets of conserved quantities that have extensive cardinality, such behaviour would manifest. A related example in the existing literature is Ref. [16] where the authors discuss the extensive entropy generation to be related to the recent developments under the rubric of “higher-form” symmetries [17]. One difference is that non-commuting conserved operators are on (non-contractible) long loops in the model of Ref. [16], whereas in the constructions discussed here, the non-commuting quantities are “local” throughout in this sense. It remains to be seen if there is a higher-form symmetry perspective on the mechanism and models discussed in this work. Another aspect is the gauge like behaviour discussed earlier. This may be a novel way in which gauge-like physical degrees of freedom emerge in quantum spin-$\frac{1}{2}$ systems, e.g. when comparing to the Levin-Wen model, Kitaev’s honeycomb model, Haah’s code and X-cube model  [18, 19, 20, 21, 22]. This mechanism can in general operate in any number of dimensions. Also may be mentioned that it is special to spin models, in particular spin-$\frac{1}{2}$, since bosons do not naturally accommodate such anticommutation and it does not appear so for fermions as well [23]. Higher spin models can accommodate more general forms of non-commutation as exemplified in Ref. [16] and local versions of non-commutation beyond anticommutation will be interesting to find. Constructing solvable models in two and higher dimensions based on the above will also be interesting.
Some physical consequences originating from this mechanism in a particular two-dimensional model not discussed here and closely related to case (a) has been discussed in Ref. [15]. Another physical point that is perhaps of relevance relates to quantum chaos bounds [24]. It has been shown that the SYK model saturates this bound [25, 26, 27]. Given the extensive ground state entropy of the SYK model [28] and the relation of zero modes to the saturation of the chaos bound [27, 29], it is tempting to conjecture that the spin models discussed here may also approach – perhaps saturate – the quantum chaos bound. These models and in particular the two-dimensional one of Fig. 3 may then provide spin models with local interactions that approach the quantum chaos bound. Finally, this work suggests a general theory for constructing models with extensive ground state entropy, say in the spirit of what Refs. [30, 31] and related papers [32, 33, 34, 35, 36, 37] do for spin models with free fermion spectra [38].
Acknowledgements:– Funding support from SERB-DST, India via Grant No. MTR/2022/000386 and partially by Grant No. CRG/2021/003024 is acknowledged.